OpenStudy (anonymous):

the manager of a music store ordered 20 new violins she ordered some of two different models-the standard and the deluxe. each standard violin costs $500, and the deluxe costs $800. if the manager spent exactly $11,500 on these violins, how many deluxe violins did she order? show your work or explain how you know you have to use the elimination process ..

OpenStudy (anonymous):

Let x = number of standard violins. Let y = number of deluxe violins. 500x + 800y = 11 500 x + y = 20 Using the process of elimination: 500x + 800y = 11 500 500x + 500y = 10 000 300y = 1500 y = 1500/300 Therefore y = 5 Therefore, 5 deluxe violins were ordered. Hope I helped

OpenStudy (anonymous):

in this part "500x + 800y = 11 500 500x + 500y = 10 000" how did you get the 10,000?

OpenStudy (anonymous):

To be able to use the process of elimination, we have to get 1 of the variables to cancel out, so the other can be isolated and solved for. Therefore, if we multiply the entire second equation by 500, we can get the x's in both equation to be equal. From there, we can just subtract both equations and isolate/solve for y.